# Can Large Language Models Act as Ensembler for Multi-GNNs?

### v1

@@ Abstract.
Graph Neural Networks (GNNs) have emerged as powerful models for learning from graph-structured data. However, GNNs lack the inherent semantic understanding capability of rich textual nodes attributes, limiting their effectiveness in applications. On the other hand, we empirically observe that for existing GNN models, no one can consistently outperforms others across diverse datasets. In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. The model first aligns multiple GNNs, mapping the representations of different GNNs into the same space. Then, through LoRA fine-tuning, it aligns the space between the GNN and the LLM, injecting graph tokens and textual information into LLMs. This allows LensGNN to ensemble multiple GNNs and leverage LLM’s strengths, leading to a deeper understanding of both textual semantic information and graph structural information. Experimental results show that LensGNN  outperforms existing models. This research advances text-attributed graph ensemble learning by providing a robust, superior solution for integrating semantic and structural information. We provide our code and data here: https://anonymous.4open.science/r/EnsemGNN-E267/.
Large Language Models, Text-attributed Graph, Graph Neural Networks
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. INTRODUCTION
@@
Graphs are structured data that captures the inter-relations between entities in the real world. To learn from graphs, graph neural networks (GNNs) have been proposed, where graph convolution is introduced (Kipf and Welling, 2016) and message passing is the main mechanism for neighborhood aggregation (Hamilton et al., 2017a; Veličković et al., 2017). GNNs have demonstrated significant success across various applications such as social network analysis (Hamilton et al., 2017a), recommendation systems (Wang et al., 2019), and molecular property prediction (Gilmer et al., 2017).
@@
Despite the success, two major challenges remain unresolved. First, GNNs, although powerful in capturing graph structures, often lack the inherent semantic understanding capability required to process the rich textual attributes of nodes (Zhao et al., 2019). This can lead to the loss of valuable semantic information during the learning process, limiting the effectiveness of GNNs in applications where node features contain meaningful texts, such as citation networks and social media platforms. Second, while there have been a series of GNNs proposed, no one has been shown to consistently lead others across datasets of various domains. For example, we compare the node classification performance of four representative GNNs, namely, APPNP (Gasteiger et al., 2018), GAT (Veličković et al., 2017), GCN (Kipf and Welling, 2016) and GIN (Xu et al., 2019) on three benchmark datasets (Yan et al., 2023): Cora, Citeseer and PubMed. For fairness, we set the equal number of hidden representation layers and also the same dimensionality size. The results given in Figure 1 demonstrate that the winner varies across the datasets. The challenge of selecting the optimal GNN for a given dataset remains unresolved, as different GNN architectures exhibit varying strengths. This also restricts the wide applicability of GNNs.
[…]
Comparison of node classification accuracy for various GNN models across different datasets. To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
Figure 1. Comparison of node classification accuracy for various GNN models across different datasets. To ensure fairness, the number of hidden layers and the size of hidden representations are kept consistent across these GNNs.
@@
To remedy the incapability of GNNs in semantic understanding, existing works (Qin et al., 2024; Li et al., 2024; Wang et al., 2024; Ye et al., 2024; Fang et al., 2024) have resorted to large language models (LLMs), which have been shown to be prominent in text understanding and generation. The inherent advantage of GNNs in utilizing graph structure, fused with the strength of LLMs in semantic understanding, mutually enhances both models and leads to superior performance in downstream tasks. However, how to select the optimal GNN in different scenarios remains a gap. Since the effectiveness of GNNs varies across datasets, there naturally arises a question: Can we develop an ensembled model for multi-GNNs? The ensembled model is expected to integrate the strengths of multiple GNNs and can consistently perform well across datasets. To further leverage the power of LLMs in text understanding, we thus upgrade the question: Can LLMs act as ensembler for multi-GNNs?
@@
In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets. The model captures semantic information from node textual attributes using LLMs, while simultaneously enabling LLMs to implicitly evaluate and combine the strengths of multiple GNNs. By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets. Specifically, since multi-GNNs could generate node embeddings in different low-dimensional spaces, we first align multi-GNNs by alternatively feeding their generated representations into a shared classifier to train them sequentially. After that, we freeze the parameters of GNNs to generate node embeddings and integrate them into the embedding layer of the used LLM, treating them as embeddings corresponding to graph tokens that encapsulate graph structural information. Subsequently, we concatenate graph tokens from multi-GNNs, text tokens from node attributes, and hand-crafted instructions to create a prompt and fine-tune the LLM using LoRA. In this way, we can not only align GNNs and LLMs, but implicitly leverage the power of LLM to ensemble multi-GNNs. Finally, the main contributions of our paper are given as follows:
  • We propose a novel method LensGNN to ensemble multi-GNNs with LLMs. To our best knowledge, we are the first work to attempt the trial. LensGNN combines the strengths of both GNNs and LLMs, which can be used to avoid the intricate process of manual GNN selection on various datasets.
  • We introduce an effective graph prompt tuning paradigm, which enhances LLMs’ capability in comprehending graph topology with two phases of alignment. The multi-GNN alignment maps multi-GNNs into the same low-dimensional space while the GNN-LLM alignment injects graph token representations into LLMs.
  • We conduct extensive experiments to show the effectiveness of LensGNN. In particular, we compare LensGNN with 17 state-of-the-art competitors. Our experimental results show that LensGNN outperforms baselines in a majority of cases, demonstrating its superior performance.
2. RELATED WORK
@@
GNNs (Kipf and Welling, 2016; Xu et al., 2019; Veličković et al., 2017; Hamilton et al., 2017b) have been widely used in learning from graph-structured data. Existing GNN studies have been conducted on various types of graphs such as heterophilic graphs (Zhu et al., 2020; Pei et al., 2020; Bi et al., 2024), signed graphs (Derr et al., 2018; Huang et al., 2019) and temporal graphs (Rossi et al., 2020; Longa et al., 2023; Xu et al., 2020). Further, to improve the effectiveness of GNNs, there are also works (Nguyen et al., 2023; Rusch et al., 2023; Black et al., 2023) devoted to solving the notorious over-smoothing (Chen et al., 2020) and over-quashing problems (Akansha, 2023). In parallel, to scale GNNs to large-scale graphs, some works (Liao et al., 2022) focus on enhancing the model efficiency. A comprehensive GNN-related survey can be found at (Wu et al., 2020). Despite their advantages, GNNs struggle to process textual attributes effectively, as traditional GNNs lack the semantic understanding capability that is necessary to handle text-attributed graphs (TAGs). In addition, to tailor a GNN model for a given dataset, existing methods mainly rely on neural architecture search (NAS) (Zoph and Le, 2017). However, it is very computationally expensive, which necessitates new exploration.
@@
Recently, LLMs have revolutionized tasks involving semantic understanding and language generation. Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies (Jin et al., 2023). One primary category of approaches takes LLMs as predictors in graph-based tasks, where models like GraphLLM (Chai et al., 2023) utilizes few-shot learning to address graph reasoning challenges. ENGINE (Zhu et al., 2024) is designed to efficiently integrate textual and topological information by adding a tunable lightweight GNN side structure alongside each layer of LLMs.Similarly, heuristic reasoning approaches, including Chain-of-Thought (CoT) (Wei et al., 2022) and models like StructGPT (Jiang et al., 2023), have demonstrated improved graph-based inference capabilities. Another notable category is to employ LLMs as encoders, as seen in the works of TextGNN (Zhu et al., 2021) and AdsGNN (Li et al., 2021), which focus on optimization strategies for encoding graph data. Employing LLMs as encoders involves using large language models to extract textual features, which are then used as initial feature vectors for nodes or edges in a graph. These feature vectors are passed into graph neural networks (GNNs) for further processing, allowing the model to incorporate both the textual information from the LLM and the structural information from the graph. Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods such as LTRN (Zhang et al., 2021) and GLEM (Zhao et al., 2022), where iterative training helps improve cross-modality learning. LLMs and GNNs are aligned through iterative pseudo-label training (prediction alignment) or by using contrastive learning to align their latent representations (latent space alignment). Finally, graph-empowered LLMs like GreaseLM (Zhang et al., 2022) and DRAGON (Yasunaga et al., 2022) illustrate the potential of enhancing language models with graph-based knowledge to enrich semantic understanding. They integrate the strengths of large language models and graph structures by modifying Transformer architectures to encode text and graph data simultaneously, allowing for a more comprehensive understanding of the interconnected information within graphs. These diverse approaches highlight the evolving landscape of LLM and GNN integration. However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
3. PRELIMINARIES
@@
(1) $h_{v}^{(t)}=\text{UPDATE}\left(\text{AGGREGATE}\left(\{h_{u}^{(t-1)}:u\in\mathcal{N}(v)\}\right),h_{v}^{(t-1)}\right),$
where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$. The AGGREGATE function is responsible for aggregating the representations of the neighboring nodes, and the UPDATE function combines the aggregated information with the previous state $h_{v}^{(t-1)}$ of node $v$ to update its current state. Through this iterative process, GNNs are able to learn increasingly rich node representations, capturing both local and global structural information in the graph. The specific implementations of the AGGREGATE and UPDATE functions can vary depending on the particular GNN model.
4. METHODOLOGY
To enable the LLM to integrate multiple GNNs, alignment is primarily required in two aspects. First, the output representations of all the GNNs need to be aligned into a common vector space. Second, there must be alignment between the output representations of the GNNs and the semantic space required by the LLM. Based on this idea, we have designed the model in the following two steps. The structure of the whole model is shown in Fig. 2.
4.1. Aligning multi-GNNs
@@
Figure 2. The overall process of LensGNN. Step 1 is the workflow of the pre-training phase of LensGNN’s GNNs. We employ a language pre-training encoder to extract semantic information from the textual attributes of nodes. Concurrently, we train the GNN encoder to capture structural information. During this phase, a MLP is utilized as a classifier, and the model is trained through gradient feedback. Step 2 is the workflow of the LoRA (Low-Rank Adaptation) supervised fine-tuning stage of LensGNN’s LLM. At this stage, we freeze the parameters of the GNNs and infer the hidden representations of the graph nodes. We segment GNN representations and convert them into graph tokens that LLM can read. The graph tokens, node text tokens, and other prompts are concatenated to form an instruction, which is then input into the LLMs for LoRA supervised fine-tuning.
[…]
In the pre-training phase, LensGNN consists of the following architecture: a pre-trained language encoder with frozen parameters, a GNN layer, a linear mapping layer, and an MLP classifier, all of which have parameters that are updated during training.In the fine-tuning phase of LensGNN, the parameters of the pre-trained language encoder and the GNN layers are frozen, and only the Lora parameters of the LLM are updated. In addition, the size of Hidden Representation is (1, t * e), where e is the length of the enbedding of each token in LLM, and t is the number of tokens each node representation is converted into.
Node feature initialization. To utilize the rich semantic information of textual features, we use pre-trained language models to initialize them. Specifically, given node $v_{i}$ with feature vector $x_{i}$, we use Sentence-BERT (Reimers, 2019) to get its initial embedding vector $\tilde{x}_{i}$ by:
(2) $\tilde{x}_{i}=\text{Sentence-BERT}\left(x_{i}\right).$
@@
(3) $H^{[k]}=\text{GNN}^{[k]}(A,\tilde{X}\mid\Theta_{G_{k}}),$
@@
Reshape node representations. The output of each GNN is then connected to a linear layer, which transforms node representations to the dimensionality of hidden embeddings in LLMs. This step paves the way for the subsequent GNN and LLM alignment. Details will be given in the next section. For each node $v_{i}$ in the $k$-th GNN, its reshaped embedding vector $\tilde{h}^{[k]}_{i}$ is denoted as
(4) $\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$
@@
4.2. Ensembling multi-GNNs with LLM
After GNNs are trained, we can get node representations capturing rich graph structural information. Although LLMs have shown competitive capability of text semantic understanding, they are ineffective in understanding graph structure (Guo et al., 2023). Therefore, in the second stage, we empower LLMs to comprehend graph representations. Further, with specially designed prompts, we enable these models to implicitly perform ensembling.
Align GNNs and LLMs. The key step in enabling LLMs to understand GNN tokens lies in aligning GNN representations, which encapsulate rich structural information, with the semantic space required by LLMs. This process necessitates fine-tuning LLMs. However, due to the extensive number of parameters in these models, fine-tuning requires significant computational resources, which is challenging for practical applications. As a result, existing approaches often freeze the parameters of LLMs and train an additional linear layer to act as a projector (Tang et al., 2024), mapping GNN representations into the semantic space. While the method makes training feasible and yields some effectiveness, the expressive capacity of the linear layer is limited, leading to the difficulty in fully leveraging the semantic understanding capabilities of LLMs.
@@ 4.2. Ensembling multi-GNNs with LLM
Figure 3. A prompt template. The prompt is divided into two parts: Instruction and Input. The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content. The latter provides the specific node text inputs and their multiple GNN representations.
@@
  1. (1)
    Simultaneous input of textual strings from a node and its adjacent neighbors, graph tokens of the node, and task specification. Existing studies (Kipf and Welling, 2016; Veličković et al., 2017; Verma et al., 2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction. Further, clear task instructions are necessary for accurate prediction.
  2. (2)
@@
Based on the above principles, we provide a prompt template used in our experiments, as shown in Figure 3. Details on the prompt are given in the caption of the figure. In practice, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously. In our experiments, we merge them into one step.
5. EXPERIMENT
In this section, we conduct extensive experiments to evaluate the performance of our model across various datasets, aiming to answer the following questions:
  • RQ1: How does LensGNN perform compared to SOTA baseline models?
  • RQ2: What is the importance of different components in LensGNN?
  • RQ3: Can small language models serve as multi-GNN ensembler?
  • RQ4: How efficient is our method LensGNN ?
  • RQ5: Is our method LensGNN insensitive to model hyper-parameters?
Table 1. Datasets statistics.
Dataset # Nodes # Edges # Classes
Cora 2,708 5,429 7
PubMed 19,717 44,338 3
ogbn-arXiv 169,343 1,166,243 40
Table 2. Hyperparameter Search Spaces.
Hyperparameter Search space
learning rate $\{2.0\times 10^{-5},5.0\times 10^{-5},1.0\times 10^{-4}\}$
dropout $[0,0.25]$
loraplus lr ratio $\{16.0,24.0,32.0\}$
Graph token $t$$\{1,2,4,8,16\}$
Table 3. The classification performance over the methods on 3 datasets. We highlight the best score on each dataset in bold and underline the runner-up’s. The - means that we did not obtain results from their original paper.

### v2

@@ Abstract
Graph Neural Networks (GNNs) have emerged as powerful models for learning from graph-structured data. However, GNNs lack the inherent semantic understanding capability of rich textual node attributes, limiting their effectiveness in applications. On the other hand, we empirically observe that for existing GNN models, no one can consistently outperforms others across diverse datasets. In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. The model first aligns multiple GNNs, mapping the representations of different GNNs into the same space. Then, through LoRA fine-tuning, it aligns the space between the GNN and the LLM, injecting graph tokens and textual information into LLMs. This allows LensGNN to ensemble multiple GNNs and take advantage of the strengths of LLM, leading to a deeper understanding of both textual semantic information and graph structural information. The experimental results show that LensGNN  outperforms existing models. This research advances text-attributed graph ensemble learning by providing a robust and superior solution for integrating semantic and structural information. We provide our code and data here: https://anonymous.4open.science/r/EnsemGNN-E267/.
Can Large Language Models Act as Ensembler for Multi-GNNs?
Hanqi Duan1, Yao Cheng1, Jianxiang Yu1, Xiang Li1,* 1East China Normal University, Correspondence: xiangli@dase.ecnu.edu.cn (Xiang Li*)
1 INTRODUCTION
@@
Graphs are structured data that captures the inter-relations between entities in the real world. To learn from graphs, graph neural networks (GNNs) have been proposed, where graph convolution is introduced Kipf and Welling (2016) and message passing is the main mechanism for neighborhood aggregation Hamilton et al. (2017a); Veličković et al. (2017). GNNs have demonstrated significant success across various applications such as social network analysis Hamilton et al. (2017a), recommendation systems Wang et al. (2019), and molecular property prediction Gilmer et al. (2017).
Figure 1: Comparison between GNN models on node classification across different datasets.
@@
Despite the success, two major challenges remain unresolved. First, GNNs, although powerful in capturing graph structures, often lack the inherent semantic understanding capability required to process the rich textual attributes of nodes Zhao et al. (2019). This can lead to the loss of valuable semantic information during the learning process, limiting the effectiveness of GNNs in applications where node features contain meaningful texts, such as citation networks and social media platforms. Second, while there have been a series of GNNs proposed, no one has been shown to consistently lead others across datasets of various domains. For example, we compare the node classification performance of four representative GNNs, namely, APPNP Gasteiger et al. (2018), GAT Veličković et al. (2017), GCN Kipf and Welling (2016) and GIN Xu et al. (2019) on three benchmark datasets Yan et al. (2023): Cora, Citeseer and PubMed. For fairness, we set the equal number of hidden representation layers and also the same dimensionality size. The results given in Figure 1 demonstrate that the winner varies across the datasets. The challenge of selecting the optimal GNN for a given dataset remains unresolved, as different GNN architectures exhibit varying strengths. This also restricts the wide applicability of GNNs.
@@
To remedy the incapability of GNNs in semantic understanding, existing works Qin et al. (2024); Li et al. (2024); Wang et al. (2024); Ye et al. (2024); Fang et al. (2024) have resorted to large language models (LLMs), which have been shown to be prominent in text understanding and generation. The inherent advantage of GNNs in utilizing graph structure, fused with the strength of LLMs in semantic understanding, mutually enhances both models and leads to superior performance in downstream tasks. However, how to select the optimal GNN in different scenarios remains a gap. Since the effectiveness of GNNs varies across datasets, there naturally arises a question: Can we develop an ensembled model for multi-GNNs? The ensembled model is expected to integrate the strengths of multiple GNNs and can consistently perform well across datasets. To further leverage the power of LLMs in text understanding, we thus upgrade the question: Can LLMs act as ensembler for multi-GNNs?
@@
In this paper, we study whether LLMs can act as an ensembler for multi-GNNs and propose the LensGNN model. Given multiple GNNs, the model integrates GNNs with LLMs, avoiding the tedious efforts in cherry-picking GNN across datasets. By dynamically integrating the outputs of different GNNs, LensGNN can not only preserve the semantic richness of node attributes, but also optimize the usage of diverse GNNs, which enhances the model’s generalizability across a wide range of tasks and datasets. Specifically, since multi-GNNs could generate node embeddings in different low-dimensional spaces, we first align multi-GNNs by alternatively feeding their generated representations into a shared classifier to train them sequentially. After that, we freeze the parameters of GNNs to generate node embeddings and integrate them into the embedding layer of the used LLM, treating them as embeddings corresponding to graph tokens that encapsulate graph structural information. Subsequently, we concatenate graph tokens from multi-GNNs, text tokens from node attributes, and hand-crafted instructions to create a prompt and fine-tune the LLM using LoRA. In this way, we can not only align GNNs and LLMs, but implicitly leverage the power of LLM to ensemble multi-GNNs. Finally, the main contributions of our paper are given as follows:
  • We propose a novel method LensGNN to ensemble multi-GNNs with LLMs.
  • We introduce an effective graph prompt tuning paradigm with both the multi-GNN alignment and the GNN-LLM alignment.
  • We conduct extensive experiments to show the superiority of LensGNN.
2 RELATED WORK
@@
GNNs Kipf and Welling (2016); Xu et al. (2019); Veličković et al. (2017); Hamilton et al. (2017b) have been widely used in learning from graph-structured data. Despite their advantages, GNNs struggle to process textual attributes effectively, as traditional GNNs lack the semantic understanding capability that is necessary to handle text-attributed graphs (TAGs). In addition, to tailor a GNN model for a given dataset, existing methods mainly rely on neural architecture search (NAS) Zoph and Le (2017). However, it is very computationally expensive, which necessitates new exploration.
@@
Recently, LLMs have revolutionized tasks involving semantic understanding and language generation. Meanwhile, the integration of LLMs and GNNs has garnered significant attention, leading to innovative methodologies Jin et al. (2023). One primary category of approaches takes LLMs as predictors in graph-based tasks, like GraphGPT Tang et al. (2024), LLaGA Chen et al. (2024) GraphLLM Chai et al. (2023) and ENGINE Zhu et al. (2024). Another notable category is to employ LLMs as encoders, as seen in the works of OFA Liu et al. (2024), TextGNN Zhu et al. (2021) and AdsGNN Li et al. (2021), which focus on optimization strategies for encoding graph data. Additionally, the alignment of LLMs and GNNs has been explored through prediction alignment methods, such as GOFA Kong et al. (2024), LTRN Zhang et al. (2021) and GLEM Zhao et al. (2022). These diverse approaches highlight the evolving landscape of LLM and GNN integration. However, there still lack studies exploring multi-GNNs ensembling with LLMs, which is particularly useful to avoid cherry-picking GNNs for a given dataset.
3 PRELIMINARIES
@@
$h_{v}^{(t)}=\text{UPDATE}\left(\text{AGG}\left(\{h_{u}^{(t-1)}:u\in\mathcal{N}(v)\}\right),h_{v}^{(t-1)}\right),$ (1)
where $\mathcal{N}(v)$ denotes the set of neighboring nodes of $v$. The AGG function is responsible for aggregating the representations of the neighboring nodes, and the UPDATE function combines the aggregated information with the previous state $h_{v}^{(t-1)}$ of node $v$ to update its current state. Through this iterative process, GNNs are able to learn increasingly rich node representations, capturing both local and global structural information in the graph. The specific implementations of the AGG and UPDATE functions can vary depending on the particular GNN model.
4 METHODOLOGY
This section introduces the main steps in LensGNN, which includes aligning multi-GNNs and ensembling multi-GNNs with LLM. The overall model procedure is given in Fig. 2.
4.1 Aligning multi-GNNs
@@
Figure 2: The overall process of LensGNN.
Node feature initialization. To utilize the rich semantic information of textual features, we use pre-trained language models to initialize them. Specifically, given node $v_{i}$ with feature vector $x_{i}$, we use Sentence-BERT Reimers (2019) to get its initial embedding vector $\tilde{x}_{i}$ by:
$\tilde{x}_{i}=\text{Sentence-BERT}\left(x_{i}\right).$ (2)
@@
$H^{[k]}=\text{GNN}^{[k]}(A,\tilde{X}\mid\Theta_{G_{k}}),$ (3)
@@
Reshape node representations. The output of each GNN is then connected to a linear layer, which transforms node representations into the dimensionality of hidden embeddings in LLMs. This step paves the way for the subsequent alignment of GNN and LLM. Details will be given in the next section. For each node $v_{i}$ in the $k$-th GNN, its reshaped embedding vector $\tilde{h}^{[k]}_{i}$ is denoted as
$\tilde{h}^{[k]}_{i}=\text{Linear}^{[k]}\left(h^{[k]}_{i}\mid\Theta_{L_{k}}\right).\$ (4)
@@
4.2 Ensembling multi-GNNs with LLM
After GNNs are trained, we can get node representations capturing rich graph structural information. Although LLMs have shown competitive capability of text semantic understanding, they are ineffective in understanding graph structure Guo et al. (2023). Therefore, in the second stage, we empower LLMs to comprehend graph representations. Further, with specially designed prompts, we enable these models to implicitly perform ensembling.
Align GNNs and LLMs. The key step in enabling LLMs to understand GNN tokens lies in aligning GNN representations, which encapsulate rich structural information, with the semantic space required by LLMs. This process necessitates fine-tuning LLMs. However, due to the extensive number of parameters in these models, fine-tuning requires significant computational resources, which is challenging for practical applications. As a result, existing approaches often freeze the parameters of LLMs and train an additional linear layer to act as a projector Tang et al. (2024), mapping GNN representations into the semantic space. While the method makes training feasible and yields some effectiveness, the expressive capacity of the linear layer is limited, leading to the difficulty in fully leveraging the semantic understanding capabilities of LLMs.
@@ 4.2 Ensembling multi-GNNs with LLM
  1. 1.
    Simultaneous input of textual strings from target node and its adjacent neighbors, graph tokens of the node, and task specification. Existing studies Kipf and Welling (2016); Veličković et al. (2017); Verma et al. (2023) have demonstrated that aggregating the information from both the center node and its adjacent neighbors could contribute to the label prediction. Further, clear task instructions are necessary for accurate prediction.
  2. 2.
@@
Based on the above principles, we provide a prompt template used in our experiments. Details on the prompt template are given in Appendix D. Further, the two steps: aligning GNNs and LLMs, and ensembing multi-GNNs can be trained simultaneously. In our experiments, we merge them into one step.
5 EXPERIMENT
Table 1: Comparison on classification accuracy (%) with MLP, GNN models and LM-based methods. LensGNN utilizes Baichuan2-13B-chat as the backbone LLM. We highlight the best score on each dataset in bold and underline the runner-up’s.